# Positive maps and extendibility hierarchies from copositive matrices

### v1

@@ 7.2. Complete extendibility
Remark 7.7.
The set of separable bosonic states are exactly equal the cone generated by product bosonic states [ISTY08], i.e
@@
If $n=2$ the question is obvious: $\mathsf{PPTBExt}_{2}^{(K_{2})}=\mathsf{PPT}(\mathbb{C}^{2}\vee\mathbb{C}^{2})=\mathsf{Sep}(\mathbb{C}^{2}\vee\mathbb{C}^{2})$. When $n\geq 3$, it is known that $\mathsf{PPT}(\mathbb{C}^{n}\vee\mathbb{C}^{n})\supsetneq\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})$ [TG09, MATS21] however explicit construction of such examples remains to be a challenging problem. Hence, it also implies that it is challenging to construct examples that are weakly entangled, and also highly extendible.
@@
Simiarly as in Proposition 7.3, we first derive a sum-of-squares correspondence to the dual of $\mathcal{K}_{r}$-PPT bosonic extendibility (i.e the set of witnesses for $\mathcal{K}_{r}$-PPT bosonic extendibility), which we prove later has a correspondence with the SOS heirarchy for LDUI/CLDUI witnesses.
Proposition 7.9.
@@
The proof is almost identical to that of Proposition 7.3, so we shall only highlight the differences rather than repeating the entire argument. The main distinction arises in the duality condition
@@
where the operators $\tilde{\rho}\in\mathcal{B}\left(\bigvee^{r}\mathbb{C}^{n}\right)^{+}$ are now constrained by $\tilde{\rho}^{\top_{[j]}}\geq 0$ for all $j=1,\ldots,\lfloor r/2\rfloor$, that is indeed the the condition that $\tilde{\rho}$ is PPT across every bipartition. Consequently, we must consider the sets
$\mathcal{C}_{0}=\mathsf{PSD}\Big(\bigvee^{r}\mathbb{C}^{n}\Big),\quad\mathcal{C}_{j}=\{X\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})^{\otimes r}:X^{\Gamma_{[j]}}\geq 0\}\;(j=1,\ldots,\lfloor r/2\rfloor),$
which together imply that $W\otimes I_{n}^{\otimes r-2}\in\big(\bigcap_{j=0}^{\lfloor r/2\rfloor}\mathcal{C}_{j}\big)^{\circ}$. The remaining arguments then follow in the same way as before. ∎
We also recall the following characterization of SOS polynomials from [ZVP06, Proposition 9].
@@
Let us consider a complex polynomial
@@
for $z\in\mathbb{C}^{n}$, where the second and third equalities follow from the duality relation Eq. 8 and the observation $z\odot\overline{z}\in\mathbb{R}^{n}$. If $X^{\mathsf{LDUI}}_{{A},{B}}\in(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$, then Proposition 7.9 implies that the real polynomial $x\in\mathbb{R}^{n}\mapsto h(x)$ is a sum of squares, and hence $A+A^{\top}+2{\rm Re}(\mathring{B})\in\mathsf{K}_{n}^{(r-2)}$.
@@
Since $\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})={\rm conv}\{|v^{\otimes 2}\rangle\langle v^{\otimes 2}|:v\in\mathbb{C}^{n}\}$ (Remark 7.7), $X^{\mathsf{LDUI}}_{{A},{B}}\in\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})^{\circ}$ if and only if
@@
The following proposition is an easy application of [PY24, Proposition 3.2], but we include its proof for completeness.
@@
Suppose $P\in(\mathsf{K}_{n}^{(r-2)})^{\circ}$ and let $W\in(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$. Then by Theorems 7.11 and 7.14, $X^{\mathsf{LDUI}}_{{A},{B}}:=\mathcal{T}_{\mathsf{LDUI}}W\in(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$ and $(A+\mathring{B})+(A+\mathring{B})^{\top}\in\mathsf{K}_{n}^{(r-2)}$. Therefore, one has
@@
The same conclusion as in Corollary 7.17 was obtained in [RPAR+25]. A careful inspection of the arguments in that paper reveals that the existence of an $r$-partite PPT Dicke state having a PPT entangled bipartite reduction is assumed without proof. Theorems 7.15 and 7.17 fills in this gap in a mathematically rigorous way.
We also refer to [ATSL12] for a construction of $n$-qubit $(n\geq 4)$ entangled states that are PPT across all bipartitions. Since we do not have the complete characterization of the hierarchy $\mathsf{PPTBExt}_{n}^{(r-1)}$, we do not know whether two hierarchies we consider are equivalent for bosonic states, or particularly mixtures of Dicke states. We pose this as the open question.
Question 7.19.
@@
The sets are equivalent to if 7.6 admits a affirmative solution, in which case the arguments in Theorem 7.15 can be applied analogously to star $r$-extendibility.
We would also like to state the following question regarding the $\mathsf{K}_{n}^{(r)}$ hierarchy and its potential implications for the theory of entanglement.
Question 7.20.
@@
For $n\geq 5$, is the inclusion $\mathsf{K}_{n}^{(r)}\subseteq\mathsf{K}_{n}^{(r+1)}$ strict for every $r\geq 1$? The affirmative answer would imply that all the inclusions in Eq. 28 are strict. Furthermore, the same strict inclusions would hold within the class of Dicke states $X^{\mathsf{LDUI}}_{{P},{P}}$.
8. Markov-Choi maps
In this section we consider an extension of Example 3.8 to the case where the first element of the pair is an arbitrary real matrix $A$ (with non-negative entries), while the second element is of the form $\operatorname{diag}(A)-\mathring{J}$. This class of examples yields $\operatorname{\mathsf{DUC}}$ maps that are very important in the literature, such as the Choi map and many of its generalizations. These maps are of the form
$\Phi^{\mathsf{DUC}}_{{A},{\operatorname{diag}(A)-\mathring{J}}}(Z)=\operatorname{diag}\{A\operatorname{diag}[Z]\}-\mathring{Z}.$
We call such maps Markov-Choi because of their precise form: on the diagonal, the map acts like a classical Markov matrix $A$, while on the off-diagonal elements, the map flips the sign of the entry, exactly like the celebrated Choi map [Cho75].
We gather some important properties of such maps in the next proposition.
Proposition 8.1.
Let $A\in\operatorname{\mathsf{EWP}}_{n}$. Then:
  1. (1)
    $(A,\operatorname{diag}(A)-\mathring{J})\in\operatorname{\mathsf{PCOP}}_{n}\iff\displaystyle{\sum_{i=1}^{n}\frac{t_{i}}{t_{i}+(At)_{i}}\leq 1}\,\quad\forall t\in\mathbb{R}^{n}_{+}$.
  2. (2)
    $(A,\operatorname{diag}(A)-\mathring{J})\in\operatorname{\mathsf{CLDUI}}^{+}_{n}\iff\displaystyle{\sum_{i=1}^{n}\frac{1}{1+A_{ii}}\leq 1}$.
  3. (3)
    $(A,\operatorname{diag}(A)-\mathring{J})\in\operatorname{\mathsf{PCP}}_{n}\iff(A,\operatorname{diag}(A)-\mathring{J})\in\mathsf{PDNN}_{n}\iff\displaystyle{\sum_{i=1}^{n}\frac{1}{1+A_{ii}}\leq 1}$ and $A_{ij}A_{ji}\geq 1$
Proof.
For the first point, let us consider the DUC map associated to the pair $(A,\operatorname{diag}(A)-\mathring{J})$:
$\Phi^{\mathsf{DUC}}_{{A},{\operatorname{diag}A-\mathring{J}}}(Z)=\operatorname{diag}\{A\operatorname{diag}[Z]\}-\mathring{J}\odot Z=\operatorname{diag}\{(I+A)\operatorname{diag}[Z]\}-Z.$
Note that the latter form of the map does not correspond to the DUC map associated to the pair (I+A,J)(I+A,-J): the two matrices in the pair do not have the same diagonal. Still, $(A,\operatorname{diag}(A)-\mathring{J})\in\operatorname{\mathsf{PCOP}}_{n}$ if and only if the corresponding map $\Phi^{\mathsf{DUC}}_{{A},{\operatorname{diag}A-\mathring{J}}}$ is positive. For an arbitrary vector $v\in\mathbb{C}^{n}$, where $v_{i}\neq 0$ for all $i\in[n]$, we have
[…]
$\iff\left\|\left(\frac{v_{i}}{\sqrt{\sum_{j=1}^{n}A_{ij}+\mathds{1}_{i=j})|v_{j}|^{2}}}\right)_{i\in[n]}\right\|\leq 1$
$\iff\sum_{i=1}^{n}\frac{|v_{i}|^{2}}{|v_{i}|^{2}+(A|v_{\cdot}|^{2})_{i}}\leq 1,$
which is the condition in the statement for $t_{i}:=|v_{i}|^{2}>0$. Conversely, the condition implies that $\Phi^{\mathsf{DUC}}_{{A},{\operatorname{diag}A-\mathring{J}}}(|v\rangle\langle v|)\geq 0$ for all $v\in\mathbb{C}^{n}$ with non-zero entries, and thus the positivity of $\Phi^{\mathsf{DUC}}_{{A},{\operatorname{diag}A-\mathring{J}}}$ by a density argument.
For the second point, using the characterization in [SN21, Section 5], the condition on the LHS is equivalent to $\operatorname{diag}(A)-\mathring{J}\in\mathsf{PSD}^{\mathbb{C}}_{n}$, which gives
$\operatorname{diag}(A)\geq\mathring{J}$ $|\displaystyle\iff\operatorname{diag}(I+A)\geq J=\outerproduct{\mathbf{1}}{\mathbf{1}}$
[…]
$\iff\left\|\left(\frac{1}{\sqrt{1+A_{ii}}}\right)_{i\in[n]}\right\|\leq 1$
$\iff\sum_{i=1}^{n}\frac{1}{1+A_{ii}}\leq 1.$
The third point follows from the comparison matrix argument from [SN21, Lemma 2.11] since $M(\operatorname{diag}(A)-\mathring{J})=\operatorname{diag}(A)-\mathring{J}$. ∎
Let us consider some simple special cases of the result above.
  • If $A=\operatorname{diag}\{a\}$ for a non-negative vector $a\in\mathbb{R}^{n}_{+}$, then $(A,\operatorname{diag}(A)-\mathring{J})\in\operatorname{\mathsf{PCOP}}_{n}$ by Proposition 3.5. On the other hand, we have
    $(A,\operatorname{diag}(A)-\mathring{J})\in\operatorname{\mathsf{CLDUI}}_{n}^{+}\iff\sum_{i}\frac{1}{1+a_{i}}\leq 1.$
  • If $A=xJ$ for some $x\geq 0$, then
    $(A,\operatorname{diag}(A)-\mathring{J})\in\operatorname{\mathsf{PCOP}}_{n}\iff x\geq 1-1/n.$
  • If $A=\operatorname{diag}\{a\}+\mathring{J}$ for some non-negative vector $a\in\mathbb{R}^{n}_{+}$, the equivalent condition in the statement reads
    $\sum_{i=1}^{n}\frac{t_{i}}{a_{i}t_{i}+\sum_{j}t_{j}}\leq\sum_{i=1}^{n}\frac{t_{i}}{\sum_{j}t_{j}}=1,$
    which is always satisfied. Hence, the pair $(\operatorname{diag}\{a\}+\mathring{J},\operatorname{diag}\{a\}-\mathring{J})$ is always $\operatorname{\mathsf{PCOP}}$. This also follows from the fact that $\operatorname{\mathsf{PCOP}}_{n}$ is a cone, Example 3.8, and the fact that $(\operatorname{diag}\{a\},\operatorname{diag}\{a\})\in\operatorname{\mathsf{PCOP}}_{n}$.
In [Ha03], the following further special case of Proposition 8.1 was discussed in great detail, the author providing an explicit condition for positivity.
Proposition 8.2.
[Ha03, Theorem 2.4] Let $a,c_{1},c_{2},\ldots,c_{n}$ be non-negative real numbers and consider the matrix
$A=\begin{bmatrix}a-1&0&\cdots&0&c_{1}\\ c_{2}&a-1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&a-1&0\\ 0&0&\cdots&c_{n}&a-1\end{bmatrix}$.
Then
$(A,\operatorname{diag}(A)-\mathring{J})\in\operatorname{\mathsf{PCOP}}_{n}\iff a\geq n-1\quad\text{and}\quad\sqrt[n]{c_{1}c_{2}\cdots c_{n}}\geq n-a.$

### v4

@@ 8.2. Complete extendibility
Indeed, the set of separable bosonic states are exactly equal the cone generated by product bosonic states [ISTY08], i.e
@@
If $n=2$ the question is obvious: $\mathsf{PPTBExt}_{2}^{(\mathcal{K}_{2})}=\mathsf{PPT}(\mathbb{C}^{2}\vee\mathbb{C}^{2})=\mathsf{Sep}(\mathbb{C}^{2}\vee\mathbb{C}^{2})$. When $n\geq 3$, it is shown that $\mathsf{PPT}(\mathbb{C}^{n}\vee\mathbb{C}^{n})\supsetneq\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})$ [TG09, MATS21]. However, explicit construction of examples beyond this, i.e., entangled bose-symmetric states having high degree of extendibility, remains to be a challenging problem.
@@
Simiarly as in Theorem 8.3, we first derive a SOS correspondence to the dual of $\mathcal{K}_{r}$-PPT bosonic extendibility (i.e the set of witnesses for $\mathcal{K}_{r}$-PPT bosonic extendibility), which we prove later has a correspondence with the SOS hierarchy for LDUI/CLDUI witnesses.
Theorem 8.9.
@@
[…]
where $\rho\in\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})}$ and $\tilde{\rho}$ is any $\mathcal{K}_{r}$-PPT bosonic extension of $\rho$. Note that $\tilde{\rho}\in\mathcal{B}\big(\bigvee^{r}\mathbb{C}^{n}\big)^{+}$ can be chosen to have a constraint that $\tilde{\rho}^{\top_{[j]}}\geq 0$ for all $j=1,\ldots,\lfloor r/2\rfloor$, that is indeed the the condition that $\tilde{\rho}$ is PPT with respect to every bipartition. We claim that there exist PSD operators $Z_{0},\ldots,Z_{\lfloor r/2\rfloor}\in\mathcal{M}_{n}(\mathbb{C})^{\otimes r}$ such that
$\Pi_{S}(W\otimes I_{n}^{\otimes r-2})\Pi_{S}=\Pi_{S}\Big(Z_{0}+\sum_{j=1}^{\lfloor r/2\rfloor}Z_{j}^{\Gamma_{[j]}}\Big)\Pi_{S},$
where $\Pi_{S}$ is a projection of $(\mathbb{C}^{n})^{\otimes r}$ onto the symmetric space $(\mathbb{C}^{n})^{\vee r}$. This will imply that
$\langle v^{\otimes 2}|W|v^{\otimes 2}\rangle\|v\|_{2}^{2(r-2)}=\langle v^{\otimes r}|\Pi_{S}(W\otimes I_{n}^{\otimes r-2})\Pi_{S}|v^{\otimes r}\rangle=\langle v^{\otimes r}|Z_{0}+Z_{1}^{\Gamma_{[1]}}+\cdots+Z_{r-1}^{\Gamma_{[\lfloor r/2\rfloor]}}|v^{\otimes r}\rangle$
for $v\in\mathbb{C}^{n}$. Note that the extension $\tilde{\rho}\in\bigcap_{j=0}^{\lfloor r/2\rfloor}\mathcal{C}_{j}$ where $\mathcal{C}_{0}=\mathsf{PSD}\big(\bigvee^{r}\mathbb{C}^{n}\big)$ and
$\mathcal{C}_{j}=\{X\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})^{\otimes r}:X^{\Gamma_{[j]}}\geq 0\}.$
Therefore, the required condition says that $W\otimes I_{n}^{\otimes{r-2}}\in\big(\bigcap_{j=0}^{\lfloor r/2\rfloor}\mathcal{C}_{j}\big)^{\circ}=\overline{\sum_{j=0}^{\lfloor r/2\rfloor}\mathcal{C}_{j}^{\circ}}$. In other words, there exists $Z_{0},\ldots,Z_{\lfloor r/2\rfloor}$ such that
$W\otimes I_{n}^{\otimes r-2}=Z_{0}+\sum_{j=1}^{\lfloor r/2\rfloor}Z_{j}^{\Gamma_{[j]}}$
where $\Pi_{S}Z_{0}\Pi_{S}\geq 0$, $Z_{j}\geq 0$ for $j\geq 1$ (since $Z_{j}^{\top_{[j]}}\in\mathcal{C}_{j}^{\circ}$), showing the claim (The closure can be ignored as we can show that the set of Hermitian polynomials of the type
$h_{Z_{0},\ldots,Z_{\lfloor r/2\rfloor}}(z)=\sum_{j=0}^{\lfloor r/2\rfloor}\langle\overline{z}^{\otimes j}z^{\otimes r-j}|Z_{j}|\overline{z}^{\otimes j}z^{\otimes r-j}\rangle,\quad z\in\mathbb{C}^{n},$
is closed. Indeed, this set is the conic hull of the set of Hermitian forms $z\mapsto|\langle\varphi|\overline{z}^{\otimes j}z^{\otimes r-j}\rangle|^{2}$ for $\varphi\in(\mathbb{C}^{n})^{\otimes r}$ and $j=0,1,\ldots,\lfloor r/2\rfloor$.)
[…]
$\langle v^{\otimes r}|\big(W\otimes I_{n}^{\otimes r-2}-(Z_{0}+\sum_{j=1}^{\lfloor r/2\rfloor}Z_{j}^{\Gamma_{[j]}})\big)|v^{\otimes r}\rangle=0$ $\quad\forall\,v\in\mathbb{C}^{n}$
[…]
It is well-known that ${\rm span}\{|v^{\otimes r}\rangle\langle v^{\otimes r}|:v\in\mathbb{C}^{n}\}=\mathcal{B}\big(\bigvee^{r}\mathbb{C}^{n}\big)$ [Chi10, Har13], and hence one has
$\Pi_{S}(W\otimes I_{n}^{\otimes r-2})\Pi_{S}=\Pi_{S}(Z_{0}+Z_{1}^{\Gamma_{[1]}}+\cdots Z_{\lfloor r/2\rfloor}^{\Gamma_{[\lfloor r/2\rfloor]}})\Pi_{S}.$
Now for $\rho\in\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})}$ and for $\tilde{\rho}$ a $\mathcal{K}_{r}$-PPT bosonic extension of $\rho$, it is straightforward that
[…]
and therefore, $W\in(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$. ∎
Next, let us recall the following characterization of SOS polynomials from [ZVP06, Proposition 9].
@@
Recall from Proposition 8.4 that a necessary condition for $X^{\mathsf{LDUI}}_{{A},{B}}\in(\mathsf{PPTBExt}_{n}^{(r-1)})^{\circ}$ is
$A+A^{\top}+2{\rm Re}(\mathring{B})\in\mathsf{K}_{n}^{(r-2)}.$
The next result shows that this is actually also sufficient when we consider the dual of $\mathcal{K}_{r}$-PPT bosonic extendibility (note that $(\mathsf{PPTBExt}_{n}^{(r-1)})^{\circ}\subseteq(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$).
@@
Let us consider a Hermitian polynomial
@@
for $z\in\mathbb{C}^{n}$, where the second and third equalities follow from the duality relation Eq. 8 and the observation $z\odot\overline{z}\in\mathbb{R}^{n}$. If $X^{\mathsf{LDUI}}_{{A},{B}}\in(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$, then Theorem 8.9 implies that the real polynomial $x\in\mathbb{R}^{n}\mapsto h(x)$ is a sum of squares, and hence $A+A^{\top}+2{\rm Re}(\mathring{B})\in\mathsf{K}_{n}^{(r-2)}$.
@@
Since $\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})={\rm conv}\{|v^{\otimes 2}\rangle\langle v^{\otimes 2}|:v\in\mathbb{C}^{n}\}$, $X^{\mathsf{LDUI}}_{{A},{B}}\in\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})^{\circ}$ if and only if
@@
Let us also record the result analogous to Theorem 8.5, which follows directly from Theorems 8.11 and 8.12.
Corollary 8.14.
If $n\geq 6$ and $M\in\operatorname{\mathsf{COP}}_{n}\setminus\bigcup_{r\geq 0}\mathsf{K}_{n}^{(r)}$, then we have for any $N\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$,
$X^{\mathsf{LDUI}}_{{N},{M-\mathring{N}}}\in\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})^{\circ}\setminus\bigcup_{r\geq 2}(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}.$
In particular, $\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})^{\circ}=\overline{\bigcup_{r\geq 2}(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}}\supsetneq\bigcup_{r\geq 2}(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$ whenever $n\geq 6$.
Now we shift our focus to the characterization of (complete graph) extendibility of Dicke states introduced in Section 2.4. First of all, the following proposition is an easy application of [PY24, Proposition 3.2], but we include its proof for completeness.
@@
Suppose $P\in(\mathsf{K}_{n}^{(r-2)})^{\circ}$ and let $W\in(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$. Then by Theorems 8.11 and 8.15, $X^{\mathsf{LDUI}}_{{A},{B}}:=\mathcal{T}_{\mathsf{LDUI}}W\in(\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})})^{\circ}$ and $(A+\mathring{B})+(A+\mathring{B})^{\top}\in\mathsf{K}_{n}^{(r-2)}$. Therefore, one has
@@
The same conclusion as in Corollary 8.18 was obtained in [RPAR+25]. A careful inspection of the arguments in that paper reveals that the existence of an $r$-partite PPT Dicke state having a PPT entangled bipartite reduction is assumed without proof. Theorems 8.16 and 8.18 fills in this gap in a mathematically rigorous way.
We also refer to [ATSL12] for a construction of $r$-qubit ($r\geq 4$) entangled states that are PPT across all bipartitions. Since we do not have the complete characterization of the hierarchy $\mathsf{PPTBExt}_{n}^{(r-1)}$, we do not know whether two hierarchies we consider are equivalent for bosonic states, or particularly mixtures of Dicke states. We pose this as the open question.
Question 8.20.
@@
The sets are equivalent to if Question 8.7 admits a affirmative solution, in which case the arguments in Theorem 8.16 can be applied analogously to star $r$-extendibility. We would also like to state the following question regarding the $\mathsf{K}_{n}^{(r)}$ hierarchy and its potential implications for the theory of entanglement.
Question 8.21.
@@
For $n\geq 5$, is the inclusion $\mathsf{K}_{n}^{(r)}\subseteq\mathsf{K}_{n}^{(r+1)}$ strict for every $r\geq 1$? The affirmative answer would imply that all the inclusions in Eq. 35 are strict. Furthermore, the same strict inclusions would hold within the class of (mixtures of) Dicke states.